OpenStudy (anonymous):

Solve for a. 4a + 5b = 11 b = 3a - 13 When I did this, I got 4, but that isn't an answer choice. Did I do something wrong?

jimthompson5910 (jim_thompson5910):

4a + 5b = 11 4a + 5(3a - 13) = 11 ... replace b with 3a - 13 4a + 15a - 65 = 11 keep going to solve for 'a'

OpenStudy (anonymous):

Mind you, one of the answer choices was "None Of These."

jimthompson5910 (jim_thompson5910):

I'm getting a = 4 as well (just worked on it)

jimthompson5910 (jim_thompson5910):

so there's possibly a typo

OpenStudy (anonymous):

I just wanted to check if I was correct X3 I have serious problems with these kinds of problems. Thank you c:

jimthompson5910 (jim_thompson5910):

definitely something to bring up with the teacher because s/he will know how to fix it

jimthompson5910 (jim_thompson5910):

and you're doing great, so keep up the good work

OpenStudy (anonymous):

Thank you. Would you mind helping me with a couple of others? After awhile I will get it, I promise! XD

jimthompson5910 (jim_thompson5910):

sure I'd love to help you out what's your question?

jimthompson5910 (jim_thompson5910):

I'm sure you will catch on quickly, so no worries

OpenStudy (anonymous):

Solve for c. 2/3c - 1/3d = 2 1/5c + 1/10d = 1 When I did it I got d = 3 and c=9/2

jimthompson5910 (jim_thompson5910):

that's a bit off, but close though

jimthompson5910 (jim_thompson5910):

the first thing you can do is get rid of all the fractions (since fractions tend to complicate things)

jimthompson5910 (jim_thompson5910):

the common denominator in the first equation is 3 so if you were to multiply EVERY term by 3, you would cancel out the fractions 2/3c - 1/3d = 2 3*(2/3c) - 3*(1/3d) = 3*2 6c/3 - 3d/3 = 6 2c - d = 6

jimthompson5910 (jim_thompson5910):

see how I did that?

OpenStudy (anonymous):

Yeah. That makes sense. You did it to keep the whole equation proportional, right?

jimthompson5910 (jim_thompson5910):

exactly, to balance everything out, I multiplied both sides (or every term) by 3

jimthompson5910 (jim_thompson5910):

how can you use this trick to simplify the second equation?

OpenStudy (anonymous):

Multiply everything by 10

OpenStudy (anonymous):

to get 2c + d = 10

jimthompson5910 (jim_thompson5910):

very good, you're a total pro at this

jimthompson5910 (jim_thompson5910):

we have gone from this system of equations 2/3c - 1/3d = 2 1/5c + 1/10d = 1 to this system (which is equivalent) 2c - d = 6 2c + d = 10

jimthompson5910 (jim_thompson5910):

now we could isolate one variable and use substitution, but we can use elimination and get the job done faster why? because we have a -d term in the first equation and a +d term in the second equation. They add to -d+d = 0d =0, so this means that adding the two equations means the 'd' terms go away. This leaves us with just 'c' and we can solve for it so add the equations to get 2c - d = 6 + 2c + d = 10 ------------ 4c + 0d = 16 So we now know that 4c = 16

OpenStudy (anonymous):

so c would be 4

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

Thank you very much x3

jimthompson5910 (jim_thompson5910):

glad to be of help